# A Nonperturbative Toolkit for Quantum Gravity

### v2

@@ 4 Tool 2: Infinite overcompleteness and topological simplification
The result (27), computed in the saddlepoint approximation of the gravitational path integral, is a coarse-grained statement. We would like to lift it to a fine-grained statement
@@
where again each term in (29) is given by the Laplace transform of the corresponding canonical term. For example, the term $\overline{Z(\beta)_{E}Z(\beta)_{E}}$ is given by the Laplace transform of $\overline{Z(\beta)Z(\beta)}$, which in the saddle point approximation is simply $\overline{Z(\beta)}\times\overline{Z(\beta)}$, and hence $\overline{Z(\beta)_{E}Z(\beta)_{E}}=\overline{Z(\beta)}_{E}\times\overline{Z(\beta)}_{E}$. Similarly $\overline{Z(\beta)\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle}=\overline{Z(\beta)}\times\overline{\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle}$ , as there are no wormholes saddles connecting the $Z(\beta)$ boundary to those of $\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle$ because the $Z(\beta)$ boundary condition contains no shell matter insertions. Hence by the Laplace transform, and using (27), $\overline{Z(\beta)_{E}\langle\beta|\prod_{\mathcal{H}_{shell,E}}|\beta\rangle}=\overline{Z(\beta)}_{E}\times\overline{Z(\beta)}_{E}$.
@@
An analogous problem was dealt with in Balasubramanian:2024yxk ; Boruch:2024kvv by considering the $\kappa\to\infty,\frac{\kappa}{e^{1/G_{N}}}\to\infty$ limit. The calculations considered in this work involve gravitational path integrals with some index loop $\overline{G^{n}_{ij}M_{ji}\cdots}$ where $M_{ji}$ is a quantity with two shell indices. This sum over saddle geometries includes fully-connected, partially-connected and disconnected contributions. Since the shell sourced by $\mathcal{O}_{i}$ can only propagate in the bulk if absorbed by $\mathcal{O}^{\dagger}_{i}$, for every index loop implied by the boundary conditions the saddlepoint geometry must have a connected component that joins the associated bra and ket. Any saddlepoint that breaks the index loop will be suppressed by $1/\kappa$. Hence in the strict $\kappa\to\infty,\frac{\kappa}{e^{1/G_{N}}}\to\infty$ limit only wormhole geometries that support maximal index loops contribute. We will refer to this as the $\kappa\to\infty$ or large $\kappa$ limit. Note that this does not mean that only the fully connected geometry contributes in this limit. If there are multiple maximal loops, each loop just has to appear in a connected component of the saddlepoint geometry.
Putting everything together, in the large $\kappa$ limit the two maximal index loops in (30) must be unbroken, leading to $\overline{(\langle\beta|\prod_{\mathcal{H}_{shell,E}}|\beta\rangle)^{2}}=\left(\overline{\langle\beta|\prod_{\mathcal{H}_{shell,E}}|\beta\rangle}\right)^{2}=\left(\overline{Z(\beta)}_{E}\right)^{2}$ within the saddlepoint approximation. This establishes (28). Note that even in the models that offer a great deal of analytic control like JT gravity, the large $\kappa$ limit was needed to extend coarse-grained statements to fine-grained ones, see Boruch:2024kvv .
@@ 4.1 Canonical path integrals
We should be careful with the $\kappa\rightarrow\infty$ limit beause of the the $\kappa^{n+1}$ factors in Eqs. 32 and 33. Formally we will consider the theory to have some very high energy cut off, rendering the dimension of $\mathcal{H}_{LR}$ and therefore $\kappa,\frac{\kappa}{e^{1/G_{N}}}$ very large but finite. After taking the $n\to-1$ limit the cutoff can be taken to infinity upon which the $\mathcal{O}(1/\kappa)$ terms vanish. This order of limits makes (33) well defined. The extension to the fine-grained statement via $\overline{(Z(\beta)-\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle)^{2}}=0$ now follows trivially, showing the fine-grained relation (31), in the path integral approximation where we sum overall saddles. This suggests the equality of the full Hilbert spaces $\mathcal{H}_{LR}=\mathcal{H}_{shell}$.
@@
In the large shell mass limit the wormhole saddles are proportional those of the Z partition function $\overline{Z(\beta)}$. In the AdS case the Z-partition function has 3 saddles: the large black hole, thermal AdS and the small black hole. Dominance between the large BH and thermal AdS saddle is determined by the Hawking-Page transition while the small BH is always subleading. Recently, Barbon:2025bbh discussed the inabilities of the small BH due to negative modes contributing to the one-loop determinant around this saddle. Such instabilities can challenge the interpretation of the shell-wormhole saddles computing coarse-grained overlaps of a positive-definite inner product. However, in the canonical ensemble the small black holes is never the dominant saddle, and after taking all the saddles into account $\overline{Z(\beta)}$ is still positive-definite. Furthermore Barbon:2025bbh showed that in a careful treatment of the microcanonical ensemble these issues can also be resolved. Hence, we will not consider such instabilities in this work.
@@ 4.1.1 Semiclassical factorisation of the trace in the large $\kappa$ limit
shown in Balasubramanian:2024yxk , without a microcanonical projection and resolvent resummation. As discussed in Sec. 2, the action of the operators $e^{-\tilde{\beta}_{L,R}H_{L,R}}$ on the states in $\mathcal{H}_{LR}$ is to append boundary cylinders of length $\tilde{\beta}_{L,R}$ respectively to the $L,R$ cuts defining state. Defining $k_{L}=e^{-\tilde{\beta}_{L}H_{L}}$ and $k_{R}=e^{-\tilde{\beta}_{R}H_{R}}$ and using (8) we obtain
@@
Repeating this argument to show $\overline{\left(Tr_{\mathcal{H}_{Shell}}(e^{-H_{L}\tilde{\beta}_{L}}e^{-H_{R}\tilde{\beta}_{R}})-Z(\tilde{\beta}_{L})\times Z(\tilde{\beta}_{R})\right)^{2}}=0$ is now straightforward because there are no saddles that interpolate between $Z$ and $Tr$ boundaries as $Z$ contains no shell insertions. This establishes the fine-grained result (34) in the sum over saddles. The simplicity of this calculation as compared to the finite $\kappa$ microcanonical methods of Balasubramanian:2024yxk will be crucial for extending the result away from the semiclassical limit in Sec. 5.2.
@@
In general terms, the micro-canonical density of states of a theory is obtained by the Laplace transform of the thermal partition function. The left hand side of (38) is by definition the thermal partition function of the shell Hilbert space. Adopting the results of Sec. 3.2.2 for the Laplace transform of the Z-partition functions on the RHS of (38), we find $dim(\mathcal{H}_{Shell,E})|=e^{\mathbf{S}_{L}(E_{L})}e^{\mathbf{S}_{R}(E_{R})}$. Therefore, as a corollary of (38) we can count the dimension of $\mathcal{H}_{shell,E}$ after taking $\kappa\to\infty$ without having to deal with the subtleties of micro-canonically projecting the shell states themselves. Furthermore, this derivation can be performed without taking $m_{s}\to\infty$ limit, which played an important role in the state-counting arguments of Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Climent:2024trz . Away from this limit shell propagation times $\Delta T_{D,i}>0$ are determined by the junction conditions. However the details turn out to be unimportant:
@@
4.2 An approximately complete basis
In the $\kappa\to\infty$ limit we can now show the shell states span the entire $\mathcal{H}_{LR}$ Hilbert space. In particular, the equality $\mathcal{H}_{shell}=\mathcal{H}_{LR}$ can be derived by showing
@@ 4.2 An approximately complete basis
The key point, as discussed in earlier sections, is that in the large shell mass limit the shell homology regions and propagation times on the disk shrink to zero, and hence the sheet grows to cover the entire disk. In this limit the only shell mass dependence is in the universal factor $\prod_{i=1}^{n+1}Z_{m_{i}}$, which will be canceled by the shell state normalisation factors. As the shell action decouples, the contribution from the bulk geometry is simply that of the disk computing the norm of a time evolved version of the $|\Psi\rangle$ state: $\overline{Z}_{\mathcal{O}_{\Psi}}((n+1)(\beta_{L}+\beta_{R})+\tilde{\beta}_{L}+\tilde{\beta}_{R})$. [^1] Here again $\overline{Z}_{\mathcal{O}_{\Psi}}((n+1)(\beta_{L}+\beta_{R})+\tilde{\beta}_{L}+\tilde{\beta}_{R})$ collectively denotes all saddles to $\overline{\langle\Psi|\Psi\rangle}((n+1)(\beta_{L}+\beta_{R})+\tilde{\beta}_{L}+\tilde{\beta}_{R})$ and for each of these saddles individually the above construction follows. The upshot is thus that given a saddle for the norm of time-evolved version of the state |$\ket{\Psi}$ given by $\overline{Z}_{\mathcal{O}_{\Psi}}((n+1)(\beta_{L}+\beta_{R})+\tilde{\beta}_{L}+\tilde{\beta}_{R})$, the saddles for (44) can automatically be constructed in the large shell mass limit, because of the universal pinching-off behavior of the shell insertions. Normalizing the shell states and putting everything together we obtain

### v3

@@ 4 Tool 2: Infinite overcompleteness and topological simplification
The result (28), computed in the saddlepoint approximation of the gravitational path integral, is a coarse-grained statement. In particular, equation (28) shows that the projector onto the shell states acts as the identity on average. In this section, we lift this result to a fine-grained statement that holds in each realization of the ensemble computed by the path integral, while remaining agnostic about the precise nature of this ensemble. We obtained equation (28) by resumming the planar topologies and performing a microcanonical projection to enable the evaluation of this sum. When extending to fine-grained equalities, however, the topological resummations become intractable. Building on Balasubramanian et al. (2025); Boruch et al. (2024a), we demonstrate that this difficulty can be overcome by resolving the identity in the limit of an infinitely overcomplete basis, in which only fully connected topologies contribute. We argue that this limit arises naturally in the canonical ensemble and removes the need for a microcanonical projection when evaluating expressions such as (18). We demonstrate the utility of these methods by re-deriving key results from Balasubramanian et al. (2024c, b); Climent et al. (2024); Balasubramanian et al. (2025) directly in the canonical ensemble and without requiring topological resummation. We then use these methods to provide a more general argument that the shell states span at a fine grained level and proivde some explicit examples. First, we continue the example from the previous section and show the fine-grained equality:
@@
where again each term in (30) is given by the Laplace transform of the corresponding canonical term. For example, the term $\overline{Z(\beta)_{E}Z(\beta)_{E}}$ is given by the Laplace transform of $\overline{Z(\beta)Z(\beta)}$, which in the saddle point approximation is simply $\overline{Z(\beta)}\times\overline{Z(\beta)}$, and hence $\overline{Z(\beta)_{E}Z(\beta)_{E}}=\overline{Z(\beta)}_{E}\times\overline{Z(\beta)}_{E}$.  [^2] For the factorization of $\overline{Z(\beta)}\times\overline{Z(\beta)}$ considered here, we focus on real and positive $\beta$, so the microcanonical saddle points obtained for $\beta\to iT$ in Saad et al. (2018) do not contribute. In this work, we study shell states inserted on a boundary of topology $\mathbb{S}^{1}_{\beta}\times\mathbb{S}^{d-1}$, with particular emphasis on the case $d=3$. However, it was shown in Cotler and Jensen (2021a) (see also Cotler and Jensen (2022, 2021b)) that in the $\mathbb{S}^{1}_{\beta}\times\mathbb{S}^{2}$ and $\mathbb{S}^{1}_{\beta}\times\mathbb{S}^{4}$ cases, the path integral for $\overline{Z(\beta)Z(\beta)}$ admits so-called constrained instanton contributions. These connected wormhole geometries correspond to saddle points subject to the constraint of either fixed length or that the two ADM energies of the boundaries are fixed to be equal, $E_{1}=E_{2}=E$. Although we work within the saddle-point approximation here, it would be interesting to investigate whether these constrained instanton contributions can also be recovered in our framework. This could be achieved by deriving the constrained junction conditions and subsequently gluing the constrained wormholes into a connected contribution to, for example, $\overline{Z(\beta)_{E}\times\langle\beta|\prod{\mathcal{H}_{\text{shell},E}}|\beta\rangle}$ in the manner depicted in Fig. 19.
Similarly $\overline{Z(\beta)\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle}=\overline{Z(\beta)}\times\overline{\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle}$ , as there are no wormholes saddles connecting the $Z(\beta)$ boundary to those of $\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle$ because the $Z(\beta)$ boundary condition contains no shell matter insertions. Hence by the Laplace transform, and using (28), $\overline{Z(\beta)_{E}\langle\beta|\prod_{\mathcal{H}_{shell,E}}|\beta\rangle}=\overline{Z(\beta)}_{E}\times\overline{Z(\beta)}_{E}$.
@@
An analogous problem was dealt with in Balasubramanian et al. (2025); Boruch et al. (2024a) by considering the $\kappa\to\infty,\frac{\kappa}{e^{1/G_{N}}}\to\infty$ limit. The calculations considered in this work involve gravitational path integrals with some index loop $\overline{G^{n}_{ij}M_{ji}\cdots}$ where $M_{ji}$ is a quantity with two shell indices. This sum over saddle geometries includes fully-connected, partially-connected and disconnected contributions. Since the shell sourced by $\mathcal{O}_{i}$ can only propagate in the bulk if absorbed by $\mathcal{O}^{\dagger}_{i}$, for every index loop implied by the boundary conditions the saddlepoint geometry must have a connected component that joins the associated bra and ket. Any saddlepoint that breaks the index loop will be suppressed by $1/\kappa$. Hence in the strict $\kappa\to\infty,\frac{\kappa}{e^{1/G_{N}}}\to\infty$ limit only wormhole geometries that support maximal index loops contribute. We will refer to this as the $\kappa\to\infty$ or large $\kappa$ limit. Note that this does not mean that only the fully connected geometry contributes in this limit. If there are multiple maximal loops, each loop just has to appear in a connected component of the saddlepoint geometry.
Putting everything together, in the large $\kappa$ limit the two maximal index loops in (31) must be unbroken, leading to $\overline{(\langle\beta|\prod_{\mathcal{H}_{shell,E}}|\beta\rangle)^{2}}=\left(\overline{\langle\beta|\prod_{\mathcal{H}_{shell,E}}|\beta\rangle}\right)^{2}=\left(\overline{Z(\beta)}_{E}\right)^{2}$ within the saddlepoint approximation. This establishes (29). Note that even in the models that offer a great deal of analytic control like JT gravity, the large $\kappa$ limit was needed to extend coarse-grained statements to fine-grained ones, see Boruch et al. (2024a).
@@ 4.1 Canonical path integrals
We should be careful with the $\kappa\rightarrow\infty$ limit beause of the the $\kappa^{n+1}$ factors in Eqs. 33 and 34. Formally we will consider the theory to have some very high energy cut off, rendering the dimension of $\mathcal{H}_{LR}$ and therefore $\kappa,\frac{\kappa}{e^{1/G_{N}}}$ very large but finite. After taking the $n\to-1$ limit the cutoff can be taken to infinity upon which the $\mathcal{O}(1/\kappa)$ terms vanish. This order of limits makes (34) well defined. The extension to the fine-grained statement via $\overline{(Z(\beta)-\langle\beta|\prod_{\mathcal{H}_{shell}}|\beta\rangle)^{2}}=0$ now follows trivially, showing the fine-grained relation (32), in the path integral approximation where we sum overall saddles. This suggests the equality of the full Hilbert spaces $\mathcal{H}_{LR}=\mathcal{H}_{shell}$.
@@
In the large shell mass limit the wormhole saddles are proportional those of the Z partition function $\overline{Z(\beta)}$. In the AdS case the Z-partition function has 3 saddles: the large black hole, thermal AdS and the small black hole. Dominance between the large BH and thermal AdS saddle is determined by the Hawking-Page transition while the small BH is always subleading. Recently, Barbon and Velasco-Aja (2025) discussed the inabilities of the small BH due to negative modes contributing to the one-loop determinant around this saddle. Such instabilities can challenge the interpretation of the shell-wormhole saddles computing coarse-grained overlaps of a positive-definite inner product. However, in the canonical ensemble the small black holes is never the dominant saddle, and after taking all the saddles into account $\overline{Z(\beta)}$ is still positive-definite. Furthermore Barbon and Velasco-Aja (2025) showed that in a careful treatment of the microcanonical ensemble these issues can also be resolved. Hence, we will not consider such instabilities in this work.
@@ 4.1.1 Semiclassical factorisation of the trace in the large $\kappa$ limit
shown in Balasubramanian et al. (2025), without a microcanonical projection and resolvent resummation. As discussed in Sec. 2, the action of the operators $e^{-\tilde{\beta}_{L,R}H_{L,R}}$ on the states in $\mathcal{H}_{LR}$ is to append boundary cylinders of length $\tilde{\beta}_{L,R}$ respectively to the $L,R$ cuts defining state. Defining $k_{L}=e^{-\tilde{\beta}_{L}H_{L}}$ and $k_{R}=e^{-\tilde{\beta}_{R}H_{R}}$ and using (8) we obtain
@@
Repeating this argument to show $\overline{\left(Tr_{\mathcal{H}_{Shell}}(e^{-H_{L}\tilde{\beta}_{L}}e^{-H_{R}\tilde{\beta}_{R}})-Z(\tilde{\beta}_{L})\times Z(\tilde{\beta}_{R})\right)^{2}}=0$ is now straightforward because there are no saddles that interpolate between ZZ and $Tr$ boundaries as ZZ contains no shell insertions. This establishes the fine-grained result (35) in the sum over saddles. The simplicity of this calculation as compared to the finite $\kappa$ microcanonical methods of Balasubramanian et al. (2025) will be crucial for extending the result away from the semiclassical limit in Sec. 5.2.
@@
In general terms, the micro-canonical density of states of a theory is obtained by the Laplace transform of the thermal partition function. The left hand side of (39) is by definition the thermal partition function of the shell Hilbert space. Adopting the results of Sec. 3.2.2 for the Laplace transform of the Z-partition functions on the RHS of (39), we find $dim(\mathcal{H}_{Shell,E})|=e^{\mathbf{S}_{L}(E_{L})}e^{\mathbf{S}_{R}(E_{R})}$. Therefore, as a corollary of (39) we can count the dimension of $\mathcal{H}_{shell,E}$ after taking $\kappa\to\infty$ without having to deal with the subtleties of micro-canonically projecting the shell states themselves. Furthermore, this derivation can be performed without taking $m_{s}\to\infty$ limit, which played an important role in the state-counting arguments of Balasubramanian et al. (2024c, b); Climent et al. (2024). Away from this limit shell propagation times $\Delta T_{D,i}>0$ are determined by the junction conditions. However the details turn out to be unimportant:
@@
4.2 A semiclassically complete basis
[…]
The equality $\mathcal{H}_{shell}=\mathcal{H}_{LR}$ can be derived by showing
@@ 4.2 A semiclassically complete basis
In this section we argue that (41) holds (in the $\kappa\to\infty$ limit) for the subset of states $|\Psi\rangle\in\mathcal{H}_{LR}$ for which the path integral $\overline{\langle\Psi|\Psi\rangle}$ has at least some saddlepoint. By extension (41) then also holds for arbitrary superpositions of these states. If one supposes that there exists some basis set for $\mathcal{H}_{LR}$ consisting of states whose norms are dominated by a saddlepoint, then our argument shows $\mathcal{H}_{shell}=\mathcal{H}_{LR}$. We are not seeking to prove this assumption here, but give some heuristic arguments in Sec. 5.1. The shell states discussed here are analogous to coherent states, which generically provide a non-orthonormal, over-complete basis for the Hilbert space.
@@
[…]
@@ Conclusion
After the completion of this work the authors learned of the works Hsin et al. (2021); Boruch et al. (2024b) in which a similar argument is used in JT gravity to show any microcanonical gravity state can be reconstructed from a certain basis of $e^{\mathbf{S}}$ states. In light of this, our work can be seen as an explicit higher dimensional realization, and extension the fine-grained level, of this argument.
4.3 Some more examples
At the beginning of this section, we demonstrated the above argument explicitly for the case in which the state under consideration is the TFD state $|\Psi\rangle=|\beta\rangle$. Two additional simple examples can be worked out straightforwardly:
Shell state
Let $|\Psi\rangle$ be some shell state created with shell operator $\mathcal{O}_{S}$ and preparation temperatures $\tilde{\beta}_{L},\tilde{\beta}_{R}$. For brevity we consider the heavy shell mass limit $m_{s}\to\infty$, although this is not required, and fix $m_{s}$ to be distinct from the masses of the shell states defining the basis. Furthermore, we could instead choose $\mathcal{O}_{S}$ to carry some global charge, whilst our basis does not. In this case $m_{s}$ may equal one of the masses in the basis. As discussed in Sec. 3.1, the norm is given by $\overline{\braket{\Psi|\Psi}}=\overline{Z}(\tilde{\beta}_{L})\times\overline{Z}(\tilde{\beta}_{R})\times Z_{m_{s}}$, with $Z_{m_{s}}$ the shell mass dependent term. The boundary conditions defining $\overline{G^{n}_{ij}\langle\Psi|i\rangle\langle j|\Psi\rangle}$ is now simply that of the pinwheel discussed in (Fig. 8) with $n+2$ boundaries. A simple adaptation of (16) yields:
$\overline{G^{n}_{ij}\langle\Psi|i\rangle\langle j|\Psi\rangle}=Z_{m_{s}}\times\frac{\overline{Z}((n+1)\beta_{L}+\tilde{\beta}_{L})\overline{Z}((n+1)\beta_{R})+\tilde{\beta}_{R})}{\overline{Z}(\beta_{L})^{n+1}\overline{Z}(\beta_{R})^{n+1}}$ (51)
and hence $\overline{\langle\Psi|\prod_{\mathcal{H}_{Shell}}|\Psi\rangle}=\lim_{n\to-1}\overline{G^{n}_{ij}\langle\Psi|i\rangle\langle j|\Psi\rangle}=\overline{Z}(\tilde{\beta}_{L})\times\overline{Z}(\tilde{\beta}_{R})\times Z_{m_{s}}=\overline{\braket{\Psi|\Psi}}$. The cross terms required for the fine grained extension can be dealt with analogously. Consider for example (48). The connected contribution to $\overline{G^{n}_{ij}\langle\Psi|i\rangle\langle j|\Psi\rangle\langle\Psi|\Psi\rangle}$ is given by a pinwheel with $n+3$ boundaries:
$\overline{G^{n}_{ij}\langle\Psi|i\rangle\langle j|\Psi\rangle\langle\Psi|\Psi\rangle}=Z_{m_{s}}^{2}\frac{\overline{Z}((n+1)\beta_{L}+2\tilde{\beta}_{L})\overline{Z}((n+1)\beta_{R})+2\tilde{\beta}_{R})}{\overline{Z}(\beta_{L})^{n+1}\overline{Z}(\beta_{R})^{n+1}}.$ (52)
Taking $n\to-1$ we obtain:
$\overline{\langle\Psi|\prod_{\mathcal{H}_{Shell}}|\Psi\rangle\langle\Psi|\Psi\rangle}=Z_{m_{s}}^{2}\overline{Z}(2\tilde{\beta}_{L})\overline{Z}(2\tilde{\beta}_{R}),$ (53)
which can be recognized as the cylinder topology connected contribution to $\overline{\braket{\Psi|\Psi}\braket{\Psi|\Psi}}$ (see Fig. 7). Again, the other cross term can be dealt with similarly. We conclude that at a fine-grained level $\langle\Psi|\Psi\rangle=\langle\Psi|\prod_{\mathcal{H}_{Shell}}|\Psi\rangle$. In the case where $\mathcal{O}_{S}$ is charged under some global symmetry, the above argument shows that this symmetry is broken by the non-perturbative wormhole contributions, analogous to Hsin et al. (2021).
Light operator
We now consider the case where the mass $m_{\Psi}$ of the operator $\mathcal{O}_{\Psi}$ is heavy enough for its two-point function to be accurately described by the geodesic approximation, yet light enough to avoid significant back-reaction on the geometry. The norm $\overline{\braket{\Psi|\Psi}}$ then takes the schematic form:
$\overline{\braket{\Psi|\Psi}}=\sum_{X}\overline{Z}_{X}(\tilde{\beta}_{L}+\tilde{\beta}_{R})\times(\sum_{i}e^{-m_{\Psi}L^{i}_{X}(\tilde{\beta}_{L},\tilde{\beta}_{R})}),$ (54)
where $\{X\}$ label the different saddles to the Z-partition function $\overline{Z}(\tilde{\beta}_{L}+\tilde{\beta}_{R})$ and $L^{i}_{X}$ denotes the length of the $i$-th bulk geodesic connecting the boundary insertion points of the operator on the background geometry $X$. These insertion points are separated on the boundary by $\tilde{\beta}_{L},\tilde{\beta}_{R}$ the $L,R$ way around. The index $i$ here denotes the sum over windings. Crucially, the geodesic lengths $L^{i}_{X}$ are simple functions of the boundary separations $\tilde{\beta}_{L},\tilde{\beta}_{R}$. See, for example, Sec. 4 of Antonini et al. (2023) and references therein for details. As $\mathcal{O}_{\Psi}$ is not heavy enough to significancy back-react, the fully connected contribution to $\overline{G^{n}_{ij}\langle\Psi|i\rangle\langle j|\Psi\rangle}$ is given by $\Psi$-matter geodesics propagating on the folded wormhole background geometry constructed in Sec. 3.2.1 above. Recall that this wormhole was obtained by gluing a single disk onto itself. Hence the $\Psi$-matter geodesics are again simply geodesics of the Z-partition saddles with a longer boundary:
$\overline{G^{n}_{ij}\langle\Psi|i\rangle\langle j|\Psi\rangle}=\sum_{X}\overline{Z}_{X}((n+1)\beta_{L}+2\tilde{\beta}_{L}+(n+1)\beta_{R}+2\tilde{\beta}_{R})\times(\sum_{i}e^{-m_{\Psi}L^{i}_{X}((n+1)\beta_{L}+2\tilde{\beta}_{L},(n+1)\beta_{R}+2\tilde{\beta}_{R})}).$ (55)
In both (54) and (55) we sum over all winding, such the geodesics summed over are in one-to-one correspondence. It then follows that the $n\to-1$ limit of (55) recovers (54). Note that if we had kept only the leading terms, subtleties about the relative sizes of $\beta_{L},\beta_{R},\tilde{\beta}_{L},\tilde{\beta}_{R}$ arize, as for example there might be some positive integer $n$ at which the $i=0$ and i=1i=-1 windings degenerate for (55) but not for (54).
The fine-grained extension is trivial, as the $\mathcal{O}_{\Psi}$ insertions are not heavy enough to stablize new connected wormhole contributions to the cross terms such as $\overline{\langle\Psi|\prod_{\mathcal{H}_{Shell}}|\Psi\rangle\langle\Psi|\Psi\rangle}$ (see Fig. 19), and hence each of the cross terms factorises. [^5] See footnote 1 for a potential subtlety. We therefore conclude that $\langle\Psi|\Psi\rangle=\langle\Psi|\prod_{\mathcal{H}_{Shell}}|\Psi\rangle$ at the fine grained level, and hence that the shell states span these states also.